OpenStudy (anonymous):

please help! The radioactive element americium-241 has a half-life of 432 years. Suppose we start with a 25-g mass of americium-241. How much will be left after 283 years? Compute the answer to three significant digits.

OpenStudy (anonymous):

if i got i right it will be left 16,812 g if halflife is 432 years so in that time you would be left with 12,5 g and 283/432 = 0,655 so 12,5*0,655 = 8,1875 so 25-8,1875 = 16,8125 g